Enough of a monoprotic weak acid is dissolved in water to produce a 0.0136M solution. The pH of the resulting solution is 2.49 . Calculate the Ka for the acid.
1
Expert's answer
2020-04-14T01:43:09-0400

In case of weak acids pH is given by "-\log[\sqrt{CK_a}]"

substituting the values

"2.49=-\log[\sqrt{1.36\times10^{-2}\times K_a}]\\3.2\times10^{-3}=\sqrt{1.36\times10^{-2}\times K_a}\\10.24\times10^{-6}=1.36\times10^{-2}\times K_a\\7.5\times10^{-4}=K_a"